OpenStudy (anonymous):

2(x-5)^2=1 Solve the equation by extracting the square root

OpenStudy (anonymous):

\[2\left( x-5 \right)^2=1,\left( x-5 \right)^2=\frac{ 1 }{ 2 }\] \[x-5=\pm \frac{ 1 }{ 2 }\] \[x=5\pm \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

find the two values of x

OpenStudy (anonymous):

correction \[x-5=\pm \sqrt{\frac{ 1 }{ 2 }}=\pm \frac{ 1 }{ \sqrt{2} },x=5\pm \frac{ 1 }{ \sqrt{2} }\]